# How Do Multilingual Language Models Remember Facts?

### v2

@@ 5 Language Information Flow
(a) Counts of examples where the patch produces the object prediction.
(b) Histogram of patches that cause the context answer to be predicted in the patch language, $\mathcal{L}_{p}(o_{c})$.
Figure 5: Prediction tokens from (a) $\{=\mathcal{L},\neq r,\neq s\}$ and (b) $\{\neq\mathcal{L},=r,\neq s\}$ for XGLM, EuroLLM, mT5.
@@
Let $h^{l}_{t}$ be the representation at layer $l$ of the last token in the input. For a given input $p$, we take $h^{l}_{t,p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass. We refer to the example $p$ as the patch example, and $c$ as the context example (see Figure 3). By patching the last token representation in each layer and measuring the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are.
@@
An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$, which we note $(\mathcal{L}(r),\mathcal{L}(s))$. We conduct three experiments where the patch and context examples share certain input characteristics. In particular, we test: (1) $\{=\mathcal{L},\neq r,\neq s\}$, where the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, where the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, where the subject is the same but the language and relation differ. (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject. In all experiments, English is used as the patch language, $\mathcal{L}_{p}=\text{en}$.
@@
Let the patch input be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the context input be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$. Without intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively. [^12] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input. To analyze patching effects, we examine both output probabilities and predicted tokens. For probabilities, we aggregate across examples by calculating the change relative to the original (unpatched) probability, as in §4. For predicted tokens, we check if they match the patch or context object, or a variant with swapped language or relation (e.g., if $\mathcal{L}_{p}(o_{c})$ is predicted in the setups (2)-(3)). We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings. [^13] E.g. to claim the predicted token is $t=\mathcal{L}_{p}(o_{c})$ we require $t\neq\mathcal{L}_{c}(o_{c})$, as languages may share spellings (e.g., “Asia” in English and Spanish). Consequently, the number of examples varies across analyses of output probabilities and token predictions. See Table 4 and 6 for the number of examples for each model and experiment.
@@
Figure 6: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},=r,\neq s\}$, for XGLM, EuroLLM, mT5.
@@
Figure 7: Relative probability change of $\mathcal{L}_{c}(o_{c})$ when patching $\{\neq\mathcal{L},\neq r,=s\}$, for XGLM, EuroLLM, mT5.
@@
We present the prediction results in Figure 5(a), and probability plots in Figure 32. The relation information is contained in the last token representation when the green curve increases, since patching at that point starts producing the object corresponding to the relation in the patch $o_{r_{p},s_{c}}$. When the green curve decreases and the blue curve increases, the object has been fully extracted and is encoded in the last token vector, as predictions from these layers correspond to $o_{r_{p},s_{p}}$ and no information is taken from the context $c$. The layers where this happens align with the peaks of the extraction rate in English (Figure 25-27). This shows that the extraction measured by the vocabulary projection represents the point at which the object is extracted and encoded in the last token representation. If the object was encoded earlier and only decoded with the extraction event,$o_{r_{p},s_{p}}$ would be predicted from patches at earlier layers.
@@
We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 7.141414The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix F.2). We observe that for decoder-only models, patching in the early layers generally hurts the model’s performance for most languages, while, patching in the middle layers either increases the probability or results in only a minimal decrease. Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$. Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive. In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero. From the attention knockout analysis, we recall that in mT5, the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers. Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the context language in mT5, whereas it proves useful in XGLM and EuroLLM.
@@
We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  7) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range. For mT5, the probability drops at the beginning but, instead of plateauing as before, it continues to decline until it reaches zero. When compared to the former experiment (Figure 7), this suggests that the relation information is encoded in the middle layers, as $\mathcal{L}_{c}(o_{c})$ decreases earlier when the patch and context have different relations.
@@
Figure 8: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$. Top XGLM, bottom EuroLLM.
@@ 6 Conclusion
In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to prior research on English recall in monolingual autoregressive LLMs. We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are present in multilingual and monolingual models. However, we also identified notable differences, including the localization of knowledge in not only one but two MLP sites; and the joint role of MLPs and attention modules during the extraction phase in multilingual models. A key contribution of our work is the first-ever investigation of language encoding during recall, achieved through patching representations. In decoder-only models, the relation flows to the last token first, followed by the language. In contrast, in mT5, both relation and language flow to the last token at similar layers. This suggests that while relation and subject representations are multilingual and enable cross-lingual object extraction, the extraction phase itself is language-specific, as language encoding precedes extraction. These findings provide new evidence to understand the factual knowledge recall in transformer LMs, and to how decoder-only LMs resolve tasks in stages. Contributing with new directions for the study of cross-lingual transfer and knowledge localization.
7 Limitations
@@
Appendix A Related Work
Factual Knowledge Recall
Petroni et al. (2019) analyzed factual knowledge in pre-trained language models, by creating the LAMA dataset, which contains natural language templates and WikiData triplets. Subsequently, Elazar et al. (2021) developed the ParaRel dataset, which contains manually paraphrases of the LAMA templates, to measure the consistency of factual knowledge.  [^15] One may object to calling the ability to solve cloze tests as knowledge. LLMs are sometimes inconsistent and cannot always justify the propositions they generate. We side with Fierro et al. (2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution. This led to further research in multilingual settings, Kassner et al. (2021); Jiang et al. (2020) translated LAMA, creating mLAMA and X-FACTR, and evaluated the factual knowledge acquired by multilingual pre-trained models, while Wang et al. (2024a) evaluated knowledge from the perspective of knowledge edits and their impact between Chinese and English (Zhu et al., 2021). Building on this, Yin et al. (2022) curated a dataset that focuses on the cultural aspect of factual knowledge. While Fierro and Søgaard (2022) studied multilingual consistency of knowledge using mParaRel, a machine-translated version of the English templates in ParaRel, where subjects and objects are translated using WikiData identifiers. In a related effort, Qi et al. (2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language. Qi et al. (2023) used the mLAMA dataset. In this paper, we used mParaRel since it is an extended version of mLAMA, as it is a translation of the paraphrases in mParaRel.
Interpretability on Factual Knowledge Recall
As discussed earlier, Meng et al. (2022) was among the first to analyze, from a mechanistic perspective, whether a model’s knowledge could be localized; while Geva et al. (2023) conducted the first study on how information flows within a model to construct the final predicted object representation. Both studies focused on the GPT models (Brown et al., 2020), using English-language data. Later, Chughtai et al. (2024) contributed by demonstrating that models arrive at an answer through several independent components, with the sum of these components yielding the correct object. Their analysis covered both GPT and Pythia models, also using English data. Subsequently, Sharma et al. (2024) applied the mechanisms identified by Meng et al. (2022) and Geva et al. (2023) to the state-space language model Mamba (Gu and Dao, 2023), focusing on facts expressed in English. Additionally, Hernandez et al. (2024) examined whether a linear transformation of the subject is sufficient to represent various relations, analyzing GPT and Llama models with English data.
Activation Patching
Ghandeharioun et al. (2024) introduced the Patchscope framework to study what an intermediate representation encodes. In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input. More closely related to our work, Dumas et al. (2024) patched the last token representation to investigate if models encode language-agnostic concepts in a word translation task. Their findings, consistent with ours and those of Foroutan et al. (2022), support the existence of language-agnostic representations. Similarly, Wang et al. (2024b) patched the last token representation to study how relations are encoded in factual English queries, finding that the last token encodes the query’s relation at a specific computational stage—aligning with our results. In this paper, we are able to draw more general conclusions given that we study a cross-lingual factual recall setup. While Dumas et al. (2024) concluded that LLMs first resolve the output language, we find that the model first resolves the query’s relation and then the language. Thus, we interpret the last token representation as a Function Vector (FV) (Todd et al., 2024), encoding the function to solve the given task: for Dumas et al. (2024) the task was translation so the function encodes the language to which translate. Finally, we are also able to show that for decoder-only models the FV composes the language onto the relation FV, which extends the findings from Wang et al. (2024b).
Multilingual Language Models
In this work, we focused on generative models and selected mT5 (Xue et al., 2021), XGLM Lin et al. (2021), and EuroLLM (Martins et al., 2024) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation. However, many other models are available. For instance, BLOOM (Le Scao et al., 2023) is a decoder-only model trained on 46 languages, although low factual recall accuracy has been reported (Qi et al., 2023). Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English. More recent models that could have been considered include mGPT (Shliazhko et al., 2022), a decoder-only model trained on 61. On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
Language Script Family SOV/SVO
English Latin Germanic SVO
Spanish Latin Romance SVO
Vietnamese Latin Austroasiatic SVO
Turkish Latin Turkic SOV
Russian Cyrillic Slavic SVO*
Ukrainian Cyrillic Slavic SVO*
Japanese Kanji Proto-Japonic SOV
Korean Korean Koreanic SOV
Hebrew Hebrew Arabic VSO
Farsi (Persian) Perso-Arabic Indo-Iranian SOV
Arabic Arabic Arabic VSO
Table 2: Languages, their scripts, families, and sentence structures (SVO: subject-verb-object, SOV: subject-object-verb, VSO: verb-subject-object, SVO*: SVO dominant but SOV is also possible).
Appendix B Experimental Setup
@@ B.2 Computational Resources
Appendix C Causal Tracing
C.1 XGLM
@@ C.2 EuroLLM
Appendix D Attention Knockout
@@
Appendix E Extraction Event
@@
Appendix F Patching
@@
F.1 Different Relation, Different Subject
@@
F.2 Same Relation, Different Subject
@@
F.3 Different Relation, Same Subject

### v3

@@ 6 Language Information Flow
(a) Number of examples where the patch produces one of the four valid objects. The green curve (middle) represents layers where the last token’s representation encodes the relation but not yet the object.
(b) Number of examples where the patch causes the output to be $\mathcal{L}_{p}(o_{c})$ (the context object in the patch language). The last token’s representation at these layers encodes the relation to extract and in which language, but not yet the object.
Figure 7: Patches per layer triggering a specific object prediction. (a) Setup $\{=\mathcal{L},\neq r,\neq s\}$; (b) $\{\neq\mathcal{L},=r,\neq s\}$.
@@
Let $h^{l}_{t}$ be the representation at layer $l$ of the last token in the input. For a given input $p$, we take $h^{l}_{t,p}$ and patch it into the forward pass of another input $c$ at the same layer; that is, we set $h^{l}_{t,c}\leftarrow h^{l}_{t,p}$, and then continue the forward pass. We refer to the example $p$ as the patch example, and $c$ as the context example ( Figure 5). By patching the last token representation in each layer and observing the model’s predicted output, we can study when the representation contains the information about relation, language, and predicted object, and how cross-lingual these representations are. In all experiments, English is used as the patch language.
@@
An input example expresses a relation $r$ and a subject $s$ in language $\mathcal{L}$, which we note $(\mathcal{L}(r),\mathcal{L}(s))$. We conduct three experiments where the $p$ and $c$ examples share certain input characteristics: (1) $\{=\mathcal{L},\neq r,\neq s\}$, the language is the same but the relation and subject differ; (2) $\{\neq\mathcal{L},=r,\neq s\}$, the relation remains the same but the language and subject differ; and (3) $\{\neq\mathcal{L},\neq r,=s\}$, the subject is the same but the language and relation differ. (1) allows us to study how the relation is encoded while controlling for the language, whereas (2)-(3) help us explore how the interaction between relation and language affects the extraction of the object from the subject. In Figure 1 and 35 we show the aggregated results of these 3 settings, with setup (1) we localize when the object is resolved (blue), with setup (2) when the relation and language are encoded (purple), and with setup (3) when the relation is encoded (green).
@@
Let the patch input be $(\mathcal{L}_{p}(r_{p}),\mathcal{L}_{p}(s_{p}))$ and the context input be $(\mathcal{L}_{c}(r_{c}),\mathcal{L}_{c}(s_{c}))$. Without intervention, the model correctly predicts $\mathcal{L}_{p}(o_{r_{p},s_{p}})$ and $\mathcal{L}_{c}(o_{r_{c},s_{c}})$ respectively. [^14] For simplicity, we may refer to the object as $\mathcal{L}_{p}(o_{p})$ or $\mathcal{L}_{c}(o_{c})$ when $r$ and $s$ are from the same input. To analyze patching effects, we examine both output probabilities and predicted tokens. For probabilities, we aggregate across examples by calculating the change relative to the original (unpatched) probability, as in §5. For predicted tokens, we check if they match the patch or context object, or a variant with swapped language or relation (e.g., if $\mathcal{L}_{p}(o_{c})$ is predicted in the setups (2)-(3)). [^15] We only consider valid patch-context pairs where $o_{r_{p},s_{c}}$ and $o_{r_{c},s_{p}}$ exist, and for predicted token analysis, we enforce distinct spellings, e.g. to claim the predicted token is $t=\mathcal{L}_{p}(o_{c})$ we require $t\neq\mathcal{L}_{c}(o_{c})$, as languages may share spellings (e.g., “Asia” in English and Spanish). Consequently, the number of examples varies across analyses of output probabilities and token predictions. See Table 4 and 6 for the number of examples for each model and experiment.
@@
Figure 8: Percentage change in the probability of the context object $\mathbb{P}(\mathcal{L}_{c}(o_{c}))$ when patching $\{\neq\mathcal{L},=r,\neq s\}$.
@@
Figure 9: Percentage change in the probability of the context object $\mathbb{P}(\mathcal{L}_{c}(o_{c}))$ when patching $\{\neq\mathcal{L},\neq r,=s\}$.
@@
We present the prediction results in Figure 7(a) and the probability plots in Figure 36. An increase in the green curve indicates that relation information becomes available in the last token representation, as patching at that layer produces $o_{r_{p},s_{c}}$—the object corresponding to the relation in the patch example. When the green curve declines and the blue curve rises, the object has been fully extracted and encoded in the last token, since predictions now yield $o_{r_{p},s_{p}}$ and are no longer influenced by $c$. These transitions align with peaks in the extraction rate for English (Figure 28-30). This confirms that extraction, as measured by the vocabulary projection, marks the point where the object is both available and encoded in the last token representation. If the object had been encoded earlier and only decoded at the extraction point, $o_{r_{p},s_{p}}$ would have been predicted from earlier-layer patches.
@@
We present the probability of the context answer token $\mathcal{L}_{c}(o_{c})$ in Figure 9.161616The probability of $\mathcal{L}_{p}(o_{p})$ and per language plots are provided in Appendix E.2. We observe that for decoder-only models, patching in the early layers generally hurts the model’s performance for most languages, while, patching in the middle layers either increases the probability or results in only a minimal decrease. Given that the relation is the same in both examples, this suggests that by these middle layers, $r$ is encoded in $h^{l}_{p}$, and in the subsequent layers the subject and language of the context are integrated to yield the final prediction, $\mathcal{L}_{p}(r)+\mathcal{L}_{c}+s_{c}=\mathcal{L}_{c}(o_{c})$. Moreover, for some languages the relation representation from the patch is better than the one constructed using the context input, as the relative probability is positive. In the case of mT5, however, the probability of $\mathcal{L}_{c}(o_{c})$ consistently decreases, plateauing in the middle layers before dropping to zero. From the attention knockout analysis, recall that in mT5 the subject is integrated into the last token representation only after layer 15, while the relation is consistently represented throughout the middle layers. Therefore, these patching results imply that the relation encoded in the last token from the patch input does not help in retrieving the correct context object in the context language in mT5, whereas it proves useful in XGLM and EuroLLM.
@@
We observe that the probability of $\mathcal{L}_{c}(o_{c})$ (Figures  9) in decoder-only models, unlike the previous experiment, decreases early for all languages (except tr and ko), plateaus around the middle layers, and then drops to zero by the mid-layer range. For mT5, the probability drops at the beginning but, instead of plateauing as before, it continues to decline until it reaches zero. When compared to the former experiment (Figure 9), this suggests that the relation information is encoded in the middle layers, as $\mathcal{L}_{c}(o_{c})$ decreases earlier when the patch and context have different relations.
@@
Figure 10: Patches that cause the prediction to be $\mathcal{L}_{c}(o_{p})$, which shows the layers where the last token representation contains the relation but not yet the language.
@@ 7 Conclusion
In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to recall in English autoregressive LLMs. We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are present in multilingual and monolingual models. However, we also identified notable differences, including the join role of late MLPs and attention modules during the extraction phase in multilingual models. A key contribution of our work is the first-ever investigation of language encoding during recall, achieved through patching representations. In decoder-only models, the relation flows to the last token first, followed by the language. In contrast, in mT5, both relation and language flow to the last token at similar layers. This suggests that while relation and subject representations are multilingual and enable cross-lingual object extraction, the extraction phase itself is language-specific, as language encoding precedes extraction. These findings provide new evidence to understand the factual knowledge recall in transformer LMs, and to how decoder-only LMs resolve tasks in stages. Contributing with new directions for the study of cross-lingual transfer and knowledge localization.
Limitations
@@
Acknowledgments
We thank our colleagues at the CoAStaL NLP group Laura Cabello, Rita Ramos, and Israfel Salazar for valuable comments on the final manuscript. DE was supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM).
Appendix A Experimental Setup
@@ A.2 Computational Resources
Appendix B Causal Tracing
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 11: Average indirect effect on logit scores, when the subject input has been corrupted and some activations are restored to their uncorrupted values. The MHSA and MLP are restored in windows of 12% consecutive layers. Top GPT2-XL (only English data).
B.1 XGLM
@@ B.2 EuroLLM
Appendix C Attention Knockout
@@
Appendix D Extraction Event
@@
Appendix E Patching
Figure 35: Aggregated patching results (§6) for XGLM. The last token representation contains the function that solves the task. This function is formed in two stages: first, the relation to extract is encoded (green), and then the language is composed into the function (purple). Finally, the function is applied to the subject, and the predicted object is resolved (blue).
@@ Appendix E Patching
E.1 Different Relation, Different Subject
@@
E.2 Same Relation, Different Subject
@@
E.3 Different Relation, Same Subject